OpenStudy (anonymous):

multiply and simplify 3x/7*5/3

OpenStudy (anonymous):

@SavannahWillett

OpenStudy (anonymous):

@vinnv226 @mathslover

OpenStudy (anonymous):

@eSpeX @Directrix

Directrix (directrix):

Is this 3x/7*5/3 supposed to be (3x/7) * (5/3) ?

OpenStudy (anonymous):

yes

Directrix (directrix):

(3x/7) * (5/3) = Multiply numerators and denominators together. What is (3x) * 5 = ? What is 7 * 3 = ? @divagirl421

OpenStudy (anonymous):

15x and 21

Directrix (directrix):

See attachment. Divide out the common factor of 3. The final answer is ? Post what you get and then I will tell you what I got. Okay? @divagirl1421